### v1

@@ I Introduction
While 5G networks have entered the commercial deployment stage, the academic community has started the exploration of envisioned 6G networks. In particular, 6G networks are expected to achieve terabits per second (Tbps) level data rates, $10^{7}/km^{2}$ connection densities, and lower than 0.1 $ms$ latency[1]. To achieve these goals, many novel communication techniques are proposed, such as reconfigurable intelligence surfaces (RISs), integrated sensing and communication, mmWave and THz networks, vehicle-to-everything networks, and so on[2]. These novel techniques have shown satisfactory performance towards 6G requirements, but the complexity of network management also significantly increased. From 3G, 4G LTE to 5G and envisioned 6G networks, telecommunication (telecom) networks have become a complicated large-scale system, including core networks, transport networks, network edge, and radio access networks[3], and the telecom-domain tasks involve network configuration and troubleshooting, product design and coding, standard specification development, performance optimization and prediction, etc. To handle such complexity, machine learning (ML) has become one of the most promising solutions, and there have been a large number of studies on artificial intelligence (AI)/ML-enabled wireless networks, e.g., reinforcement learning-based network management[4], deep neural network-enabled channel state information (CSI) prediction[5], and federated learning for distributed model training in wireless environments[6]. For example, convex optimization has been applied to optimize network performance, but it requires problem-specific transformation for convexity. By contrast, reinforcement learning will transform the problem into a unified Markov decision process (MDP), and then interact with the environment to explore optimal policies. Compared with conventional optimization algorithms[2], reinforcement learning overcomes the complexity of dedicated problem reformulation, and can better handle environmental uncertainties, e.g., the growing diversity of user preferences, and more distributed and heterogeneous resources in future telecom networks. These studies have demonstrated the importance of incorporating ML to improve the efficiency, reliability, and quality of telecom services.
@@
Recently, large language model (LLM) techniques have attracted considerable interest from both academia and industry. Unlike previous ML algorithms, these large-scale models with a huge amount of parameters have shown versatile comprehension and reasoning capabilities in various fields such as health care[7], law[8], finance[9], education, and so on [10]. For instance, Wu et al. introduced a BloombergGPT model that is trained on a wide range of financial data with 50 billion parameters, and the Med-PaLM2 developed by Google achieves 86.5% correct rate on the medical question answering dataset[7]. Existing studies have demonstrated that LLM technologies have many promising features such as in-context learning (ICL), step-by-step learning, and instruction following[11]. For instance, ICL means that the model can learn from natural language instruction or task demonstration [12], and then generate expected output rapidly without any extra model training. Such an efficient learning approach can be attractive for the telecom domain, since many wireless applications require rapid or even real-time responses.
@@
Despite the great potential, LLM’s real-world application is still at a very early stage, especially for domain-specific scenarios. For instance, telecom is a broad field that includes various knowledge domains, e.g., signal transmissions, protocols, network architectures, devices, and different standards. LLM is expected to properly understand and generate content that aligns with real-world details and specific requirements of telecom applications[13]. However, such specific telecom-related requirements are rare in the existing knowledge base of general-domain LLMs. Therefore, applying a general-domain LLM directly to telecom tasks may lead to poor performance. Meanwhile, fine-tuning LLM models on telecom datasets may improve LLM’s performance of domain-specific tasks, but the telecom-specific dataset collection and filtering still require careful design and evaluation. In addition, many telecom tasks require multi-step planning and thinking, e.g., a simple coding task can include multiple steps, indicating dedicated prompting and analyses from the telecom perspective[14].
@@
Given the above opportunities and challenges, this work presents a comprehensive survey of LLM-enabled telecom networks. Different from existing studies that focus on one specific aspect such as edge intelligence [15, 16], grounding and alignment[17], this work provides a comprehensive survey on fundamentals, key techniques, and applications of LLM-enabled telecom. To be specific, this work focuses on generative models that were originally developed for language tasks, i.e., language models, and it also involves more diverse techniques and broad application scenarios such as optimization and prediction problems. Although LLM development is originally motivated by natural language tasks, it is worth noting that there have been diverse state-of-the-art explorations that are beyond the conventional language processing tasks, e.g., coding and debugging [18], recommendation [19], LLM-enabled agents [20], instruction-based optimization [21], network time-series prediction and decision making [22], etc. These LLM-inspired techniques have become crucial pillars of LLM studies, and exploring these techniques is crucial to take full advantage of LLM capabilities. Fig.1 presents the organization of this work, in which the left side indicates the telecom scenarios and demand, and the right side shows the LLM-enabled techniques111 On the high level, the terms “LLM” and “foundation model” can be used interchangeably. In this survey, the term ”foundation models” refers to models specifically developed from scratch for applications that are beyond pure language-related tasks, such as the prediction foundation models in Section VII-B, while ”LLM-enabled” or ”LLM-aided” approaches denote methods that repurpose existing pre-trained language models for telecommunication tasks.. To better present the detailed application scenarios, the bottom of Fig.1 shows telecom environments that include radio access networks, network edge, central cloud, and other network elements such as regular users, malicious users, mmWave beam, environment image sensing, RISs, backhaul traffic, etc. In particular, we focus on the following topics:
@@
1) LLM fundamentals: Understanding LLM fundamentals is the prerequisite for developing advanced applications in telecom networks[10]. Compared with existing studies[15, 16, 17], this work presents a more comprehensive overview of the model architecture, pre-training, fine-tuning, inference and utilization, and evaluation. Additionally, it presents different approaches to deploy LLMs in telecom networks, such as central cloud, network edge, and mobile LLM [16]. It further analyzes LLM fundamentals from the telecom application perspective, e.g., training or fine-tuning telecom-specific LLM models, and the importance of prompting and multi-step planning techniques for telecom tasks.
@@
3) LLM-based classification for telecom: Classification is a common task in the telecom field, and we present LLM-enabled network attack classification and detection, telecom text, image, and traffic classification problems. For instance, there have been many studies on visioned-aided blockage prediction and beamforming for 6G networks [25], and some LLM can provide zero-shoot image classification capabilities, overcoming the training difficulties of conventional algorithms in complicated signal transmission environments [26].
@@
5) LLM-aided prediction in telecom: Prediction techniques are crucial for telecom networks, such as CSI prediction[5], prediction-based beamforming[25], and traffic load prediction[32]. Existing studies have started exploring one-model-for-all time-series models. After pre-training on a large corpus of diverse time-series data, such a model can learn the hidden temporal patterns, and then generalize well across different prediction tasks without extra training. We will first introduce how to pre-train foundation models, and then present frozen pre-trained and fine-tune-based LLM models. In addition, the potential of multi-modal LLM is discussed for telecom prediction tasks.
6) Challenges and future directions: Finally, we identify the challenges and future directions of LLM-empowered telecom. The challenges focus on telecom-domain LLM model training, practical LLM deployment, and prompt engineering for telecom applications. The future directions include LLM-enabled planning, model compression and fast inference, overcoming hallucination problems, retrieval augmented-LLM, and economic and affordable LLM models.
In summary, the main contribution of this work is that we provide a comprehensive survey of the principles, key techniques, and applications for LLM-enabled telecom networks, ranging from LLM fundamentals to novel LLM-inspired generation, classification, optimization and prediction techniques along with telecom applications. This work covers nearly 20 telecom application scenarios and LLM-inspired novel techniques, aiming to be a roadmap for researchers to use LLMs to solve various telecom tasks. The rest of this paper is organized as Fig. 1. Section II discusses related surveys, and Section III presents LLM fundamentals. Section IV, V, VI, and VII focus on generation, classification, optimization, and prediction problems and telecom applications, respectively. Finally, Section VIII identifies the challenges and future directions, and Section IX concludes this work.
@@ II Related Surveys
Table I compares this work with existing studies [15, 16, 17, 28, 29, 30, 31], including LLM fundamental techniques such as pre-training and fine-tuning, and other key topics ranging from question answering to multi-modality. Firstly, Table I shows that most existing studies focus on the fundamental techniques of LLM models, e.g., pre-training LLM models for telecom tasks in general[15, 16]. Meanwhile, LLM deployment is also discussed in many existing studies, including central cloud [15, 31], network edge [16], and mobile execution[17]. Due to the storage and computational resources constraint at the network edge, Lin et al. also summarized various techniques in [16] that can be used to improve the LLM training efficiency at the network edge, such as parameter-efficient fine-tuning, split edge learning, and quantized training. In addition, sensing has become an important part of future 6G networks, and the multi-modal LLM are discussed in several existing studies, e.g., integrated sensing and communication with LLM[17, 29], multi-modal input to LLM models for intelligent sensing and communication[30], and multi-modal sensing[31]. These studies are very valuable explorations of LLM-enabled telecom networks by focusing on model training and deployment. However, LLM techniques are rapidly progressing and many LLM-inspired novel techniques and applications have been recently proposed. This work is different from existing studies in the following aspects:
@@
TABLE II: Summary of existing general and domain-specific LLM models.
@@
Moreover, Table II summarizes some existing LLM models, including general-domain LLM such as GPT-4 and Llama-2, and domain-specific LLM for healthcare[38], finance[9], time series, autonomous driving[39], and recommendation system[40]. Table II demonstrates that LLM techniques have received considerable attention in many fields, and researchers have trained various domain-specific LLM models for their application scenarios. Therefore, given the rapid progress and great potential of LLM technologies, it calls for a comprehensive investigation of how to apply LLM to the telecom field. To this end, this work answers such a question: What are the most state-of-the-art techniques inspired by LLM models, and how these techniques can be used to solve telecom domain problems? The answer to this question is crucial for building intelligent next-generation telecom networks in the 6G era.
@@ III LLM Fundamentals
This Section will introduce LLM fundamentals, and the overall organization is shown in Fig.2. It presents a thorough overview of LLM fundamentals, including the model architecture, pre-training, fine-tuning, inference and utilization, and model evaluation. We further discuss how LLM models can be deployed in telecom networks such as central cloud, network edge, and mobile devices. Finally, we analyze LLM fundamentals from the telecom application perspective, e.g., training or fine-tuning LLM models for the telecom domain.
@@ III-A Model Architecture
The fundamental component of contemporary LLM models is the transformer scheme [41], which leverages an attention mechanism to capture global dependencies between inputs and outputs. Transformers process raw inputs by tokenizing them and applying embeddings and positional encodings. The vanilla transformer architecture comprises two main components: the encoder and the decoder. The encoder’s role is to extract features and understand the relationships among all input tokens. It employs self-attention, also referred to as bidirectional attention, allowing each token to attend to every other token in both directions. Conversely, the decoder is responsible for producing the output sequence while taking into account the input sequence and previously generated tokens. It initially applies a masked attention mechanism, known as causal attention, ensuring that the current token only attends to previously generated tokens. Additionally, the decoder employs cross-attention, where the query comes from the decoder, and the key and value are from the encoder, enabling the decoder to integrate information from both the input sequence and the already generated tokens. Beyond the basic version of the attention mechanism, various variants are developed to capture the different relationships among tokens, such as multi-head attention [41], multi-query attention [42], and grouped-query attention [43]. Current architectures can be classified into three distinct categories: encoder-only architecture, encoder-decoder architecture, and decoder-only architecture.
@@ III-A3 Decoder-only architecture
Decoder-only architectures specialize in unidirectional attention, allowing each output token to attend only to its past tokens and itself. Both prefix and output tokens undergo identical processing within the decoder. Decoders are further distinguished based on their attention mechanisms into causal and non-causal decoders. In causal decoders, every token is restricted to attending to its past tokens and itself; in non-causal decoders, prefix tokens can attend to all tokens within the prefix. Causal decoders are predominantly adopted in popular LLM models, such as the GPT series [12], PaLM [49], and LLaMA [50]. Non-causal decoders [51] resemble encoder-decoder frameworks in their ability to bidirectionally process the prefix sequence and autoregressively generate output tokens sequentially.

### v2

@@ I Introduction
While 5G networks have entered the commercial deployment stage, the academic community has started the exploration of envisioned 6G networks. In particular, 6G networks are expected to achieve terabits per second (Tbps) level data rates, $10^{7}/km^{2}$ connection densities, and lower than 0.1 $ms$ latency[1]. To achieve these goals, the International Telecommunication Union (ITU) has defined six key use cases for envisioned 6G networks[2]. Specifically, three cases are extensions of IMT-2020 (5G), namely immersive communication, hyper-reliable and low-latency communication, and massive communication, and the other three novel usage cases are ubiquitous connectivity, integrated sensing and communication, and AI and communication. These novel techniques have shown satisfactory performance towards 6G requirements, but the complexity of network management also significantly increased. From 3G, 4G LTE to 5G and envisioned 6G networks, telecommunication (telecom) networks have become a complicated large-scale system, including core networks, transport networks, network edge, and radio access networks[3]. Moreover, 6G ubiquitous connectivity aims to address presently uncovered areas, e.g., rural and sparsely populated areas, by integrating other access systems such as satellite communications. In addition, 6G integrated sensing and communication is designed to improve applications requiring sensing capabilities, i.e., assisted navigation, activity detection, and environmental monitoring. Despite the potential benefits, such highly integrated network architecture and functions may lead to a huge burden on 6G network management, including network configuration and troubleshooting, product design and coding, standard specification development, performance optimization and prediction, etc.
@@
To handle such complexity, machine learning (ML) has become one of the most promising solutions, and there have been a large number of studies on artificial intelligence (AI)/ML-enabled wireless networks, e.g., reinforcement learning-based network management[4], deep neural network-enabled channel state information (CSI) prediction[5], and federated learning for distributed model training in wireless environments[6]. For example, convex optimization has been applied to optimize network performance, but it requires problem-specific transformation for convexity. By contrast, reinforcement learning will transform the problem into a unified Markov decision process (MDP), and then interact with the environment to explore optimal policies. Compared with conventional optimization algorithms[7], reinforcement learning overcomes the complexity of dedicated problem reformulation, and can better handle environmental uncertainties, e.g., the growing diversity of user preferences, and more distributed and heterogeneous resources in future telecom networks. These studies have demonstrated the importance of incorporating ML to improve the efficiency, reliability, and quality of telecom services.
@@
Recently, large language model (LLM) techniques have attracted considerable interest from both academia and industry. Unlike previous ML algorithms, these large-scale models with a huge amount of parameters have shown versatile comprehension and reasoning capabilities in various fields such as health care[8], law[9], finance[10], education, and so on [11]. For instance, Wu et al. introduced a BloombergGPT model that is trained on a wide range of financial data with 50 billion parameters, and the Med-PaLM2 developed by Google achieves 86.5% correct rate on the medical question answering dataset [8]. LLM technologies have many promising features such as in-context learning (ICL), step-by-step learning, and instruction following[12]. Existing studies have shown that LLMs can answer telecom-domain questions, generate troubleshooting reports, develop project code, and configure networks, which will significantly lower the difficulty of 6G ubiquitous connectivity management. Meanwhile, for 6G integrated sensing and communication, LLMs can understand and process multi-modal data, e.g., text, satellite or street camera images, 3D LiDAR maps and videos. It provides a promising approach to simulate and understand the 3D wireless signal transmission environment.
@@
Despite the great potential, LLM’s real-world application is still at a very early stage, especially for domain-specific scenarios. For instance, telecom is a broad field that includes various knowledge domains, e.g., signal transmissions, protocols, network architectures, devices, and different standards. LLM is expected to properly understand and generate content that aligns with real-world details and specific requirements of telecom applications[13]. However, such specific telecom-related requirements are rare in the existing knowledge base of general-domain LLMs. Therefore, applying a general-domain LLM directly to telecom tasks may lead to poor performance. Meanwhile, fine-tuning LLMs on telecom datasets may improve LLM’s performance of domain-specific tasks, but the telecom-specific dataset collection and filtering still require careful design and evaluation. In addition, many telecom tasks require multi-step planning and thinking, e.g., a simple coding task can include multiple steps, indicating dedicated prompting and analyses from the telecom perspective[14].
@@
Given the above opportunities and challenges, this work presents a comprehensive survey of LLM-enabled telecom networks. Different from existing studies that focus on one specific aspect such as edge intelligence [15, 16], grounding and alignment[17], this work provides a comprehensive survey on fundamentals, key techniques, and applications of LLM-enabled telecom. To be specific, this work focuses on generative models that were originally developed for language tasks, i.e., language models, and it also involves more diverse techniques and broad application scenarios such as optimization and prediction problems. In this survey, the term “foundation models” refers to models specifically developed from scratch for applications that are beyond pure language-related tasks, such as the prediction foundation models in Section VII-B, while “LLM-enabled” or “LLM-aided” approaches denote methods that repurpose existing pre-trained language models for telecom tasks. Moreover, when referring to LLMs, it means that the inputs to the model are purely text, and the model generates purely text as outputs, even if the model can accept inputs in other modalities, such as GPT-4V and GPT-4o. When discussing the multi-modal inputs, we explicitly describe them as multi-modal large language models or multi-modal LLMs.
Although LLM development is originally motivated by natural language tasks, it is worth noting that there have been diverse state-of-the-art explorations that are beyond the conventional language processing tasks, e.g., coding and debugging [18], recommendation [19], LLM-enabled agents [20], instruction-based optimization [21], network time-series prediction and decision making [22], etc. These LLM-inspired techniques have become crucial pillars of LLM studies, and exploring these techniques is crucial to take full advantage of LLM capabilities. Fig.1 presents the organization of this work, in which the left side indicates the telecom scenarios and demand, and the right side shows the LLM-enabled techniques. To better present the detailed application scenarios, the bottom of Fig.1 shows telecom environments that include radio access networks, network edge, central cloud, and other network elements such as regular users, malicious users, mmWave beam, environment image sensing, RISs, backhaul traffic, etc. Meanwhile, we categorize key telecom applications into generation, classification, optimization, and prediction problems to better distinguish different scenarios and customized designs [^1] Note that although the classification, optimization, and prediction capabilities are all based on the LLM’s inference and generation capabilities, this organization can significantly reduce the reader’s difficulty in understanding the LLM’s potential for telecom applications.
In particular, we focus on the following topics:
@@
1) LLM fundamentals: Understanding LLM fundamentals is the prerequisite for developing advanced applications in telecom networks[11]. Compared with existing studies[15, 16, 17], this work presents a more comprehensive overview of the model architecture, pre-training, fine-tuning, inference and utilization, and evaluation. Additionally, it presents different approaches to deploy LLMs in telecom networks, such as central cloud, network edge, and mobile LLM [16]. It further analyzes LLM fundamentals from the telecom application perspective, e.g., training or fine-tuning telecom-specific LLMs, and the importance of prompting and multi-step planning techniques for telecom tasks.
@@
3) LLM-based classification for telecom: Classification is a common task in the telecom field, and we present LLM-enabled network attack classification and detection, telecom text, image, and traffic classification problems. For instance, there have been many studies on visioned-aided blockage prediction and beamforming for 6G networks [25], and some LLM can provide zero-shot image classification capabilities, overcoming the training difficulties of conventional algorithms in complicated signal transmission environments [26].
@@
5) LLM-aided prediction in telecom: Prediction techniques are crucial for telecom networks, such as CSI prediction[5], prediction-based beamforming[25], and traffic load prediction[36]. Existing studies have started exploring one-model-for-all time-series models. After pre-training on a large corpus of diverse time-series data, such a model can learn the hidden temporal patterns, and then generalize well across different prediction tasks without extra training. We will first introduce how to pre-train foundation models, and then present frozen pre-trained and fine-tune-based LLMs. In addition, the potential of multi-modal LLM is discussed for telecom prediction tasks.
6) Challenges and future directions: Finally, we identify the challenges and future directions of LLM-empowered telecom. The challenges focus on telecom-domain LLM training, practical LLM deployment, and prompt engineering for telecom applications. The future directions include LLM-enabled planning, model compression and fast inference, overcoming hallucination problems, retrieval augmented-LLM, and economic and affordable LLMs.
In summary, the main contribution of this work is that we provide a comprehensive survey of the principles, key techniques, and applications for LLM-enabled telecom networks, ranging from LLM fundamentals to novel LLM-inspired generation, classification, optimization and prediction techniques along with telecom applications. This work covers nearly 20 telecom application scenarios and LLM-inspired novel techniques, aiming to be a roadmap for researchers to use LLMs to solve various telecom tasks. The rest of this paper is organized as Fig. 1. Section II discusses related surveys, and Section III presents LLM fundamentals. Sections IV, V, VI, and VII focus on generation, classification, optimization, and prediction problems and telecom applications, respectively. Finally, Section VIII identifies the challenges and future directions, and Section IX concludes this work.
@@ II Related Surveys
Table I compares this work with existing studies [15, 16, 17, 28, 29, 30, 31, 32, 33, 34, 35], including LLM fundamental techniques such as pre-training and fine-tuning, and other key topics ranging from question answering to multi-modality. Firstly, Table I shows that most existing studies focus on the fundamental techniques of LLMs, e.g., pre-training LLMs for telecom tasks in general [15, 16, 32, 33, 34, 35]. LLM deployment is discussed in many existing studies, including central cloud [15, 31], network edge [16], and mobile execution[17]. Due to the storage and computational resources constraint at the network edge, Lin et al. also summarized various techniques in [16] that can be used to improve the LLM training efficiency at the network edge, such as parameter-efficient fine-tuning, split edge learning, and quantized training.
Meanwhile, researchers have investigated various network application scenarios for LLM and generative AI (GAI), such as integrated satellite-aerial-terrestrial networks [32], secure physical layer communication [37], semantic communication [38], and vehicular networks [39]. For instance, Javaid et al. studied the application of LLMs to integrated satellite-aerial-terrestrial networks, including resource allocation, traffic routing, network optimization, etc [32]. Huang et al. presents a general overview of LLM for networking, involving network design, configuration, and security [35]. Du et al. present a novel concept named “AI-generated everything”, discussing the interactions between (GAI) and different network layers [40]. In addition, sensing has become an important part of future 6G networks, and the multi-modal LLM are discussed in several existing studies, e.g., integrated sensing and communication with LLM[17, 29], multi-modal input to LLMs for intelligent sensing and communication[30], and multi-modal sensing[31]. These studies are very valuable explorations of LLM-enabled telecom networks by focusing on model training and deployment. However, LLM techniques are rapidly progressing and many LLM-inspired novel techniques and applications have been recently proposed. This work is different from existing studies in the following aspects:
@@
TABLE II: Summary of existing general and domain-specific LLMs.
@@
Moreover, Table II summarizes various general and domain-specific LLMs, demonstrating that LLMs have received considerable attention across many fields. Researchers have trained various domain-specific LLMs for their application scenarios, including healthcare[46], finance[10], time series[47], autonomous driving[48], and recommendation systems[49], etc. For instance, DriveGPT4 is designed to provide interpretable end-to-end autonomous driving[48]. LLM is also used for the automated design of reward functions in robot control[44], achieving better performance than human manual designs. Thus, given the rapid progress and great potential of LLMs, a comprehensive survey is expected to summarize the latest and potential applications of LLMs in the telecom field. To this end, this work answers such a question: What are the most state-of-the-art techniques inspired by LLMs, and how can these techniques be used to solve telecom domain problems? The answer to this question is crucial for building intelligent next-generation telecom networks.
@@ III LLM Fundamentals
This Section will introduce LLM fundamentals, and the overall organization is shown in Fig.2. It presents a thorough overview of LLM fundamentals, including the model architecture, pre-training, fine-tuning, inference and utilization, and model evaluation. We further discuss how LLMs can be deployed in telecom networks such as central cloud, network edge, and mobile devices. Finally, we analyze LLM fundamentals from the telecom application perspective, e.g., training or fine-tuning LLMs for the telecom domain.
@@ III-A Model Architecture
The fundamental component of contemporary LLMs is the transformer scheme [50], which leverages an attention mechanism to capture global dependencies between inputs and outputs. Transformers process raw inputs by tokenizing them and applying embeddings and positional encodings. The vanilla transformer architecture comprises two main components: the encoder and the decoder. The encoder’s role is to extract features and understand the relationships among all input tokens. It employs self-attention, also referred to as bidirectional attention, allowing each token to attend to every other token in both directions. Conversely, the decoder is responsible for producing the output sequence while taking into account the input sequence and previously generated tokens. It initially applies a masked attention mechanism, known as causal attention, ensuring that the current token only attends to previously generated tokens. Additionally, the decoder employs cross-attention, where the query comes from the decoder, and the key and value are from the encoder, enabling the decoder to integrate information from both the input sequence and the already generated tokens. Beyond the basic version of the attention mechanism, various variants are developed to capture the different relationships among tokens, such as multi-head attention [50], multi-query attention [51], and grouped-query attention [52]. Current architectures can be classified into three distinct categories: encoder-only architecture, encoder-decoder architecture, and decoder-only architecture.
@@ III-A3 Decoder-only architecture
Decoder-only architectures specialize in unidirectional attention, allowing each output token to attend only to its past tokens and itself. Both prefix and output tokens undergo identical processing within the decoder. Decoders are further distinguished based on their attention mechanisms into causal and non-causal decoders. In causal decoders, every token is restricted to attending to its past tokens and itself; in non-causal decoders, prefix tokens can attend to all tokens within the prefix. Causal decoders are predominantly adopted in popular LLMs, such as the GPT series [58], PaLM [59], and LLaMA [60]. Non-causal decoders [61] resemble encoder-decoder frameworks in their ability to bidirectionally process the prefix sequence and autoregressively generate output tokens sequentially.
Note that LLM is a complicated system, and there are multiple approaches to apply LLMs to the telecom field, ranging from pre-training and fine-tuning, to prompting. For instance, pre-training an LLM from scratch by using telecom-domain datasets, fine-tuning a general domain LLM for specific telecom tasks, or using general domain LLMs by prompting directly. The following will introduce the key procedures and features of each approach, in which Sections III-B, III-C, and III-D introduce the procedures of pre-training, fine-tuning, and prompting, respectively.